# MicroDreamer: Zero-shot 3D Generation in $\sim$20 Seconds by Score-based Iterative Reconstruction

### v2

@@ Abstract
Optimization-based approaches, such as score distillation sampling (SDS), show promise in zero-shot 3D generation but suffer from low efficiency, primarily due to the high number of function evaluations (NFEs) required for each sample. In this paper, we introduce score-based iterative reconstruction (SIR), an efficient and general algorithm mimicking a differentiable 3D reconstruction process to reduce the NFEs. Given a single set of images sampled from a multi-view score-based diffusion model, SIR repeatedly optimizes 3D parameters, unlike the single-step optimization in SDS. With other improvements in training, we present an efficient approach called MicroDreamer that generally applies to various 3D representations and 3D generation tasks. In particular, retaining a comparable performance, MicroDreamer is 5-20 times faster than SDS in generating neural radiance field and takes about 20 seconds to generate meshes from 3D Gaussian splatting on a single A100 GPU, halving the time of the fastest zero-shot baseline, DreamGaussian. Our code is available at https://github.com/ML-GSAI/MicroDreamer.
Figure 1: MicroDreamer can generate high-quality meshes in about 20 seconds on a single A100, built on a multi-view diffusion model without additional 3D data.
1 Introduction
@@
Recently, optimization-based approaches [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13] particularly score distillation sampling (SDS) [1, 2] have emerged as promising avenues for 3D generation based on text-to-image diffusion models [14, 15, 16, 17, 18, 19]. These approaches are appealing due to their minimal, or even zero reliance on 3D data, in contrast to the data-intensive requirements of feed-forward approaches [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. This advantage is particularly significant given that 3D data are costly and scarce. Despite their promising capabilities, optimization-based approaches suffer from low efficiency due to the extensive number of function evaluations (NFEs), i.e. forward passes of the diffusion model, required for each 3D object generation.
@@
In comparison, the multi-step reconstruction process of 3D representations that enable differentiable rendering, such as neural radiance field (NeRF) [35, 36] and 3D Gaussian splatting (3DGS) [37], produce 3D contents extremely fast because they do not involve large generative neural networks. However, such approaches rely on true 3D data (i.e. abundant real multi-view images like 100 in NeRF [35]), making it unfeasible to text-to-3D and image-to-3D generation tasks. There exist works [38, 39, 40, 41] for 3D generation attempt to generate multi-views first to reconstruct 3D object directly, while such methods may suffer from texture ambiguity [39] in generated views and are limited to specific 3D representations with a long reconstruction time (see Sec. 5 for comparison).
@@
This paper introduces an efficient and general algorithm termed score-based iterative reconstruction (SIR) for zero-shot 3D generation inspiring from reconstruction and iterative optimization to address the computational issues in Sec. 3. Similar to SDS, SIR iteratively updates 3D parameters, leveraging a multi-view diffusion model without additional 3D data. However, in each iteration, SIR distinguishes itself by repeatedly optimizing 3D parameters given a set of images produced by the diffusion model, mimicking the efficient 3D reconstruction process to reduce the total NFEs (see Fig. 4). To obtain 3D-consistent and high-quality images as the ground truth for better reconstruction in each iteration, we carefully design a hybrid forward process and a sampling process to refine the images rendered from the 3D object optimized in the latest iteration.
@@
We delve into critical components beyond the algorithm to ensure compatibility with SIR. The comprehensive system is named by MicroDreamer, highlighting its efficiency for 3D generation. As detailed in Sec. 4, we introduce an initialization strategy for 3D objects, an annealed time schedule for diffusion, additional losses on reference images for image-to-3D, and a refinement procedure for deriving high-quality meshes from 3DGS.
@@
SIR and MicroDreamer are broadly applicable to both NeRF and 3DGS and both text-to-3D and image-to-3D tasks, as detailed in Sec. 5. Employing three widely adopted multi-view diffusion models [42, 43, 44], we systematically compare SIR and SDS for NeRF generation. Retaining a comparable performance, SIR can accelerate the generation process by 5 to 20 times. Besides, MicroDreamer can efficiently produce 3DGS and further refine it into high-quality meshes, delivering consistent 3D meshes in about 20 seconds on a single A100 GPU – about twice as fast as the most competitive optimization-based alternatives, DreamGaussian [9]. Remarkably, MicroDreamer is on par with speed compared to feed-forward methods trained on extensive 3D data, with a very competitive CLIP similarity [45, 9].
@@
Figure 2: Overview of SIR. SIR is a zero-shot 3D generation method that marries the strengths of reconstruction and iterative optimization. SIR reutilizes the samples from diffusion multiple times through reconstruction, reducing the total NFEs and improving efficiency.
2 Background
@@
We present background on 3D representations, diffusion models, multi-view diffusion models, and current optimization-based 3D generation methods.
2.1 3D representation
@@
Neural radiance fields [35, 36] (NeRF) and 3D Gaussian splatting [37] (3DGS) have emerged as popular 3D representations. NeRF employs an MLP to predict the colors and density of the input space coordinates. 3DGS consists of multiple 3D Gaussians parameterized by the colors, center, scale, and rotation quaternion. We denote the corresponding tunable parameters in both representations as $\theta$. Given camera poses $c$, both approaches define a differentiable rendering process, denoted by $g(\theta,c)$. They are proven efficient and effective in 3D reconstruction [35, 37] and generation [1, 9].
2.2 Diffusion models
@@
$x_{t}=\alpha_{t}x_{0}+\sigma_{t}\epsilon:=\textrm{Noise-adding}(x_{0},0\rightarrow t),\quad\epsilon\in\mathcal{N}(0,I),$ (1)
@@
Instead of using Eq. (2) directly, the sampling process of diffusion gradually denoises through the noise prediction network and generates images from pure Gaussian noise. Among existing samplers [46, 47, 48, 49, 50, 51], the denoising diffusion implicit models (DDIM)  [52] facilitate a sequence of samplers with random noise control $\eta$. When $\eta=0$, it solves the equivalent probability flow ordinary differential equation (ODE) [16] of the diffusion model and enjoys a fast sampling process with fewer number of function evaluations (NFEs) [^1] Throughout the paper, we refer to the number of forward passes through $\epsilon_{\phi}$ as NFEs.. In this setting, the one-step sampling of DDIM is given by:
$x_{t-1}=\frac{\alpha_{t-1}}{\alpha_{t}}x_{t}+\left(\sigma_{t-1}-\frac{\alpha_{t-1}}{\alpha_{t}}\sigma_{t}\right)\epsilon_{\phi}(x_{t},t):={\textrm{Sampler}}(x_{t},t\rightarrow t-1).$ (3)
@@
$x_{t+1}=\frac{\alpha_{t+1}}{\alpha_{t}}x_{t}+\left(\sigma_{t+1}-\frac{\alpha_{t+1}}{\alpha_{t}}\sigma_{t}\right)\epsilon_{\phi}(x_{t},t):={\textrm{Inversion}}(x_{t},t\rightarrow t+1).$ (4)
2.3 Multi-view diffusion models
After being trained on a modest amount of 3D data, diffusion models can generate 3D-consistent multi-view, known as multi-view diffusion models. Among them, MVDream [42] takes text inputs and outputs multi-view images consistent in 3D. In contrast, Zero-1-to-3 [6] and ImageDream [44] focus on the Image-to-3D task, taking an additional reference image as input. These models output new viewpoint images consistent with the reference image. This paper directly utilizes these pre-trained multi-view diffusion models without any further fine-tuning. For more details, please refer to the original papers and the source code in our supplementary materials. Notably, such multi-view diffusion models cannot provide a sufficient number of consistent multi-view images for 3D reconstruction directly, as shown in Fig. 7(a). Given that the multi-view diffusion in [38, 39, 40, 41] often incorporate with additional 3D priors and do not align with our setting in Sec. 4, we choose not to employ them in this paper.
2.4 Optimization-based algorithms for 3D generation
Built upon (multi-view) diffusion models, optimization-based methods (see Sec. 6 for a review) aim to generate 3D content in a zero-shot manner. Among them, score distillation sampling (SDS) [1, 2] is the most representative and popular approach. Formally, denoting the rendered images as $x=g(\theta,c)$, SDS repeats adding noise to $x$ according to Eq. (1) and updating the 3D parameters $\theta$ by
$\nabla_{\theta}\mathcal{J}_{\textrm{SDS}}(x=g(\theta);\phi):=\mathbb{E}_{t}\left[w(t)(\epsilon_{\phi}(\alpha_{t}x+\sigma_{t}\epsilon,t)-\epsilon)\frac{\partial x}{\partial\theta}\right],$ (5)
where $w(t)$ is a fixed weighting function and we omit the dependency on the prompt $y$ and camera $c$ for simplicity. Notably, by reparameterization [4] according to Eq. (2), SDS has an equivalent data-prediction form:
@@ 2.4 Optimization-based algorithms for 3D generation
3 Score-based iterative reconstruction
We present the general framework of score-based iterative reconstruction (SIR), an efficient and versatile algorithm combining both the advantages of differentiable 3D reconstruction and optimization methods in Sec. 3.1. SIR capitalizes on the efficiency of 3D reconstruction and aims to minimize the number of function evaluations (NFEs) typically required by existing methods while ensuring comparable quality through iterations. For better reconstruction, we introduce a carefully designed diffusion-based process to produce refined multi-view images as ground truth in Sec. 3.2.
3.1 Mimicking 3D reconstruction to reduce the NFEs
Existing zero-shot 3D generation algorithms [1, 2, 4, 8] iteratively optimize the 3D parameters based on a 2D diffusion (see Eq. (5) for SDS as an example). Consequently, these methods necessitate high NFEs, presenting a critical challenge to computational efficiency.
Naturally, reutilizing the outcomes of diffusion for successive updates of the 3D object—mimicking the process of differentiable 3D reconstruction could substantially decrease the overall NFEs required. However, 3D reconstruction typically requires a sufficient number of consistently aligned multi-view images, which cannot be directly obtained from the current multi-view diffusion models [6, 42, 44, 43] (see Sec. 6 for a review and discussion). Therefore, similar to existing optimization-based methods, we introduce a reconstruction-based algorithm with multiple iterations of optimizing dubbed score-based iterative reconstruction (SIR), detailed as follows.
Formally, SIR consists of $K$ reconstruction iterations. In the $k$-th iteration, where $k=0,1,\ldots,K-1$, given initial 3D parameters $\theta_{0}^{(k)}$, we randomly select several camera poses $c^{(k)}$ (detailed in Sec. 4) and employ rendering function $g(\cdot,\cdot)$ to obtain multi-view images of the current 3D object as follows:

### v3

# MicroDreamer: Efficient 3D Generation in $\sim$20 Seconds by Score-based Iterative Reconstruction

@@ Abstract
Optimization-based approaches, such as score distillation sampling (SDS), show promise in zero-shot 3D generation but suffer from low efficiency, primarily due to the high number of function evaluations (NFEs) required for each sample and the limitation of optimization confined to latent space. This paper introduces score-based iterative reconstruction (SIR), an efficient and general algorithm mimicking a differentiable 3D reconstruction process to reduce the NFEs and enable optimization in pixel space. Given a single set of images sampled from a multi-view score-based diffusion model, SIR repeatedly optimizes 3D parameters, unlike the single-step optimization in SDS. With other improvements in training, we present an efficient approach called MicroDreamer that generally applies to various 3D representations and 3D generation tasks. In particular, MicroDreamer is 5-20 times faster than SDS in generating neural radiance field while retaining a comparable performance and takes about 20 seconds to create meshes from 3D Gaussian splatting on a single A100 GPU, halving the time of the fastest optimization-based baseline DreamGaussian with significantly superior performance compared to the measurement standard deviation. Our code is available at https://github.com/ML-GSAI/MicroDreamer.
Index Terms:
3D Generation, Diffusion Model, Multi-view Diffusion, Score Distillation Sampling
I Introduction
@@
Recently, optimization-based approaches [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] particularly score distillation sampling (SDS) [1, 2] have emerged as promising avenues for 3D generation based on text-to-image diffusion models [20, 21, 22, 23, 24, 25]. These approaches are appealing due to their minimal, or even zero reliance on 3D data, in contrast to the data-intensive requirements of feed-forward approaches [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53]. This advantage is particularly significant given that 3D data are costly and scarce. Despite their promising capabilities, optimization-based approaches suffer from low efficiency due to the extensive number of function evaluations (NFEs), i.e. forward passes of the diffusion model, required for each 3D object generation. Moreover, when adopting the latent diffusion model (LDM) [23] framework, most of these approaches can only compute loss in latent space rather than pixel space, which requires backpropagation through an encoder [54, 55] and further lowers the efficiency. Even if the loss function can be written in a data-predicting form [4], this type of loss is challenging to optimize effectively (see evidence in Fig. 11). The fastest approach DreamGaussian [8] still requires about 40 seconds to generate a 3D object even employing 3D Gaussian splatting [56].
Figure 1: MicroDreamer surpasses the fastest optimization-based baseline DreamGaussian [8] in terms of both efficiency and sample quality. The optimization-based methods are highlighted in red. See Tab. III for a comprehensive comparison with more baselines.
Figure 2: MicroDreamer can generate a high-quality mesh (as illustrated above) in about 20 seconds on a single A100, built on a multi-view diffusion model without additional 3D data. See supplementary materials for 3D visualization.
@@
In comparison, the multi-step reconstruction process of 3D representations that enable differentiable rendering, such as neural radiance field (NeRF) [57, 58] and 3D Gaussian splatting (3DGS) [56], produces 3D contents extremely fast because they do not involve large generative neural networks. However, such approaches rely on true 3D data, i.e. abundant real multi-view images, making them unfeasible for text-to-3D and image-to-3D generation tasks. There exist works [59, 60, 61, 62, 63, 64, 65, 66] for 3D generation attempt to generate multi-views first to reconstruct 3D object directly. Still, such methods may require the diffusion model to simultaneously generate multi-view images with their corresponding 3D priors [60, 63], and they require longer than one minute to reconstruct a 3D object. (see Sec. II for a review and Sec. VI for comparison).
@@
To speed up the generation process, this paper presents an efficient and general 3D generation algorithm termed score-based iterative reconstruction (SIR), leveraging reconstruction to reduce NFEs and enable optimization in pixel space. Like SDS, SIR iteratively updates 3D parameters, leveraging a multi-view diffusion model without additional 3D data or 3D prior. However, in each iteration, SIR distinguishes itself by repeatedly optimizing 3D parameters given a set of images produced by the diffusion model, mimicking the efficient 3D reconstruction process to reduce the total NFEs (see Fig. 7). To obtain 3D-consistent and high-quality images as the ground truth for better reconstruction in each iteration, we carefully design a hybrid forward process and a sampling process to refine the images rendered from the 3D object optimized in the latest iteration. Besides, even mapped back to pixel space through the decoder in LDM, the refined images are still of high quality for reconstruction, enabling optimization in pixel space to speed up further (see Fig. 10).
@@
We provide a general and compatible configuration for SIR, and the comprehensive system is named MicroDreamer, highlighting its efficiency for 3D generation. As detailed in Sec. V, we introduce an initialization strategy for 3D objects, an annealed time schedule for diffusion, additional losses on reference images for image-to-3D, and a refinement procedure for deriving high-quality meshes from 3DGS.
@@
Comprehensive studies demonstrate the generality and efficiency of our proposed method. In particular, SIR and MicroDreamer broadly apply to NeRF and 3DGS and both text-to-3D and image-to-3D tasks, as detailed in Sec. VI. Employing three widely adopted multi-view diffusion models [67, 68, 69], we systematically compare SIR and SDS for NeRF generation. Retaining a comparable performance, SIR can accelerate the generation process by 5 to 20 times. Besides, MicroDreamer can efficiently produce 3DGS and further refine it into high-quality meshes, delivering consistent 3D meshes in about 20 seconds on a single A100 GPU – about twice as fast as the most competitive optimization-based alternatives, DreamGaussian [8], with significantly (compared to the measurement standard deviation) superior performance. Remarkably, MicroDreamer is on par with speed compared to representative feed-forward methods [70] trained on extensive 3D data, with a very competitive CLIP similarity [71].
@@
Figure 3: Overview of SIR. SIR is an optimization-based 3D generation method that marries the strengths of reconstruction and iterative optimization. SIR reutilizes the samples from diffusion multiple times through reconstruction, reducing the total NFEs, enabling optimization in pixel space, and improving efficiency.
II Related work
Optimization-based 3D generation. Built upon text-to-image diffusion models, optimization-based approaches [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] usually generate 3D objects without additional 3D data. Among them, the most relevant line of work [1, 7, 9, 4, 2, 17, 18, 19] proposes various distillation algorithms. Besides SDS [1], ProlificDreamer [7] proposes variational score distillation (VSD) to produce high-fidelity 3D objects via variational inference. LucidDreamer [9] incorporates DDIM inversion to strengthen the forward process. ReconFusion [15] utilizes diffusion priors to generate novel views for single-step optimization in sparse view reconstruction tasks, aiming to enhance the quality of scene reconstruction. Though promising, many of these methods suffer from inefficiency problems due to large NFEs and optimization in latent space. Distinct from existing algorithms, SIR mimics 3D reconstruction processes to optimize the 3D parameters multiple times given a set of images produced by diffusion. Our experiments demonstrate that SIR is 5-20 times faster than SDS [1] on NeRF and twice as fast as the most competitive baseline DreamGaussian [8] with better generation quality on meshes.
Feed-forward 3D generation. Contrary to optimization-based methods, feed-forward methods [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53] use large-scale 3D datasets [72, 73] training to achieve a process that can directly generate 3D objects. Early methods are characterized by their speed but often produce lower-quality 3D structures with simple textures. Several recent methods [70, 32, 36, 37, 48, 53] have exhibited the feasibility of training a Transformer utilizing more data to achieve reliable 3D content from a single image. Generating 3D content from a single view can be quite challenging. Therefore, several follow-up works [38, 39, 40, 49, 51, 42, 43, 45] have leveraged multi-view diffusion models to generate multiple views as input and then trained a feed-forward model to produce 3D content from these views. Remarkably, MicroDreamer is on par with speed compared to such methods trained on extensive 3D data, with a very competitive 3D quality measured by CLIP similarity [71].
Multi-view prediction based 3D generation. There is also a line of work [59, 60, 61, 62, 63, 64, 65, 66] dedicated to enhancing the output of multi-view diffusion models by training on 3D datasets to reconstruct 3D objects using a single reconstruction process with no or few iterations. Some approaches like Wonder3D [60] typically rely on 3D prior knowledge and are limited to specific 3D representations with a long reconstruction time. Another aspect of works like IM-3D [64] and V3D [65] involves fine-tuning the video diffusion model using 3D data and employing the generated 3D-aware multi-view images as ground truth for reconstruction. The efficiency of these methods is limited by the sampling efficiency of the video diffusion model and the long reconstruction time. Consequently, the total time required to generate a 3D object using these methods typically exceeds one minute. In contrast, thanks to the carefully designed iterative process in SIR, MicroDreamer is more efficient and applies to various 3D representations (see Tab. III for comparison).
III Background
@@
We present background on 3D representations, diffusion models, multi-view diffusion models, and current optimization-based 3D generation methods sequentially.
III-A 3D representation
@@
Neural radiance fields [57, 58] (NeRF) and 3D Gaussian splatting [56] (3DGS) have emerged as popular 3D representations. NeRF employs an MLP to predict the colors and density of the input space coordinates. 3DGS consists of multiple 3D Gaussians parameterized by the colors, centers, scales, and rotation quaternions. We denote the corresponding tunable parameters in both representations as $\theta$. Given camera poses $c$, both approaches define a differentiable rendering process, denoted by $g(\theta,c)$. They are proven efficient and effective in 3D reconstruction [57, 56] and generation [1, 8].
III-B Diffusion models
@@
$x_{t}$ $=\alpha_{t}x_{0}+\sigma_{t}\epsilon$
[…]
@@
Instead of using Eq. (2) directly, the sampling process of diffusion gradually denoises through the noise prediction network and generates images from pure Gaussian noise. Among existing samplers [74, 75, 76, 77, 78, 79], the denoising diffusion implicit models (DDIM) [80] facilitate a sequence of samplers with random noise control $\eta$. When $\eta=0$, it solves the equivalent probability flow ordinary differential equation (ODE) [22] of the diffusion model and enjoys a fast sampling process with a small number of function evaluations (NFEs) [^1] Throughout the paper, we refer to the number of forward passes through $\epsilon_{\phi}$ as NFEs.. In this setting, the one-step sampling of DDIM is given by:
$x_{t-1}$ $=\frac{\alpha_{t-1}}{\alpha_{t}}x_{t}+\left(\sigma_{t-1}-\frac{\alpha_{t-1}}{\alpha_{t}}\sigma_{t}\right)\epsilon_{\phi}(x_{t},t)$
$:={\textrm{Sampler}}(x_{t},t\rightarrow t-1).$ (3)
@@
$x_{t+1}$ $=\frac{\alpha_{t+1}}{\alpha_{t}}x_{t}+\left(\sigma_{t+1}-\frac{\alpha_{t+1}}{\alpha_{t}}\sigma_{t}\right)\epsilon_{\phi}(x_{t},t)$
$:={\textrm{Inversion}}(x_{t},t\rightarrow t+1).$ (4)
III-C Multi-view diffusion models
After being trained on a modest amount of 3D data, diffusion models can generate 3D-consistent multi-view, known as multi-view diffusion models. Among them, MVDream [67] takes text inputs and outputs multi-view images consistent in 3D. In contrast, Zero-1-to-3 [6] and ImageDream [69] focus on the Image-to-3D task, taking an additional reference image as input. These models output new viewpoint images consistent with the reference image. This paper directly utilizes these pre-trained multi-view diffusion models without any further fine-tuning. Notably, such multi-view diffusion models cannot provide sufficient consistent multi-view images for 3D reconstruction directly (see Fig. 12(a)). Given that the multi-view diffusion models in [59, 60, 61, 62, 64, 63] often incorporate additional 3D priors and do not align with our setting in Sec. V, we choose not to employ them in this paper.
III-D Optimization-based algorithms for 3D generation
Built upon (multi-view) diffusion models, optimization-based methods (see Sec. II for a review) aim to generate 3D content in a zero-shot manner. Among them, score distillation sampling (SDS) [1, 2] is the most representative and popular approach. Formally, denoting the rendered images as $x=g(\theta,c)$, SDS repeats adding noise to $x$ according to Eq. (1) and updating the 3D parameters $\theta$ by
$\nabla_{\theta}\mathcal{J}_{\textrm{SDS}}(x=g(\theta);\phi)$
::=:= $\mathbb{E}_{t}\left[w(t)(\epsilon_{\phi}(\alpha_{t}x+\sigma_{t}\epsilon,t)-\epsilon)\frac{\partial x}{\partial\theta}\right],$ (5)
where $w(t)$ is a fixed weighting function, we omit the dependency on the prompt $y$ and camera $c$ for simplicity. Notably, by reparameterization [4] according to Eq. (2), SDS has an equivalent data-prediction form:
@@ III-D Optimization-based algorithms for 3D generation
IV Score-based iterative reconstruction
We first analyze factors contributing to the efficiency bottleneck of existing work in Sec. IV-A, motivating score-based iterative reconstruction (SIR), an efficient and versatile algorithm combining the advantages of differentiable 3D reconstruction and optimization methods in Sec. IV-B. We introduce a carefully designed diffusion-based process to produce refined multi-view images as ground truth for better reconstruction in Sec. IV-C. Besides, we demonstrate how SIR facilitates the optimization of 3D content directly within pixel space in Sec. IV-D.
IV-A Efficiency bottleneck of existing work
Figure 4: Time proportion in SDS optimization. We record the time proportions of all components in SDS on 3DGS. The bottleneck lies in the large NFEs and updating in latent space.
As a motivation of SIR, we analyze the efficiency bottleneck of existing optimization-based methods [1, 2, 4, 9, 17, 18, 7]. We take the widely adopted SDS [1] as a representative example and our analysis also applies to other algorithms.
On the one hand, according to Eq. (5), SDS iteratively optimizes the 3D parameters based on a 2D diffusion. It necessitates high NFEs because it requires a forward pass of the 2D diffusion in each update of the 3D parameters. On the other hand, when employed within the latent diffusion model (LDM) [23] upon a variational auto-encoder (VAE) [83, 55], SDS computes the loss in latent space. This process requires backpropagation through the VAE encoder, further lowering generation efficiency. These arguments are validated by the empirical results in Fig. 4 quantitatively.
As for the latent space optimization, we emphasize that although SDS has a corresponding data-prediction form in Eq. (6), it is nontrivial to map the predicted $x_{0}^{(t)}$ back to pixel space through the VAE decoder for loss calculation. This is because the single-step prediction in Eq. (2) yields poor samples in the pixel space, leading to suboptimal 3D results (See experiments in Tab. 11 of Sec. VI).
IV-B Mimicking 3D reconstruction to reduce the NFEs
Motivated by the analysis of SDS, we propose SIR to reduce NFEs and enable optimization in pixel space to enhance efficiency. Naturally, reutilizing the outcomes of diffusion for successive updates of the 3D object—mimicking the process of differentiable 3D reconstruction could substantially decrease the overall NFEs required and enable optimization in pixel space. However, 3D reconstruction typically requires a sufficient number of consistently aligned multi-view images, which cannot be directly obtained from the current multi-view diffusion models [6, 67, 69, 68] (see Sec. II for a review and discussion). Therefore, similar to existing optimization-based methods, we introduce a reconstruction-based algorithm with multiple iterations of optimizing dubbed score-based iterative reconstruction (SIR), detailed as follows.
Formally, SIR consists of $K$ reconstruction iterations. In the $k$-th iteration, where $k=0,1,\ldots,K-1$, given initial 3D parameters $\theta_{0}^{(k)}$, we randomly select several camera poses $c^{(k)}$ (detailed in Sec. V) and employ rendering function $g(\cdot,\cdot)$ to obtain multi-view images of the current 3D object as follows: